OpenStudy (anonymous):

Solve 6^x = 24 and the answer is not 4

OpenStudy (experimentx):

ln(24)/ln(6)

OpenStudy (anonymous):

6^x=24 ln 6^x = ln 24 x ln 6 =ln 24 x=(ln24)/(ln6) =1.77

OpenStudy (anonymous):

yes thanks you

OpenStudy (anonymous):

You are welcome :D